OpenStudy (rational):

\[1/1, 3/2, 7/5, 17/12, 41/29, ?\]

OpenStudy (anonymous):

what is the question?

OpenStudy (freckles):

I see some fibinocci stuff going on

OpenStudy (freckles):

or however you spell it

OpenStudy (freckles):

1,1,2,3,5,8, oops nevermind

OpenStudy (freckles):

I thought I see some adding of previous terms in there

OpenStudy (rational):

interesting... thats possible i guess but my knowledge in this topic limited so not so sure how to link this sequence to fibonacci..

OpenStudy (freckles):

1/1 3/2 1+1 , 1+2 7/5 3+2, 5+2 17/12 7+5, 5+12 41/29 17+12, 12+29

OpenStudy (freckles):

so we have 41+29, 29+(41+29)

OpenStudy (freckles):

\[\frac{29+(41+29)}{41+29}\]

OpenStudy (freckles):

one sec checking my pattern

OpenStudy (freckles):

the next bottom should be 41+29 the next top should be 29+(new bottom)

OpenStudy (rational):

that works nicely xD

OpenStudy (irishboy123):

\[\frac{1}{1},\frac{3}{2},\frac{7}{5},\frac{17}{12},\frac{41}{29},\frac{99}{70},\]

OpenStudy (rational):

that looks simple compared to the method i am using! : \[a_n = 1 + \dfrac{1}{2+\dfrac{1}{2+\dfrac{1}{2+\cdots\text{(n times)}} }}\]

OpenStudy (rational):

\[a_1 = 1 = 1/1\] \[a_2 = 1 = 1 + \dfrac{1}{2} = \dfrac{3}{2}\] \[a_3=1+\dfrac{1}{2+\dfrac{1}{2}} = 1+\dfrac{2}{5} = \dfrac{7}{5}\] \[\cdots \]

OpenStudy (rational):

easy to see.. keep doing this forever gives us the pythaogrean number \(\sqrt{2}\)

OpenStudy (freckles):

I was eating sorry I'm trying currently to write my thingy in like a general form It kinda seems like it might be tricky

OpenStudy (rational):

getting a recurrence relation is easy, but an explicit relation is like finding a formula for nth digit of of sqrt(2) right ?

OpenStudy (freckles):

like how do you say this in a pretty way: \[\text{ term }=\frac{ \text{ previous numerator } +2 \text{ previous denominator }}{ \text{ previous numerator }+ \text{ previous denominator }}\]

OpenStudy (rational):

\(a_n = \dfrac{p_n}{q_n};~~p_1 = 1, q_1 = 1,~p_2=3,~q_2=2\) \(p_n = 2*p_{n-1} + p_{n-2}\) \(q_n = 2*q_{n-1} + q_{n-2}\)

OpenStudy (rational):

not completely sure if that works... still checking.. .

OpenStudy (freckles):

well p3 is 7 and q3 is 5 so an is 7/5

OpenStudy (freckles):

oops a3*

OpenStudy (freckles):

seems like it is working

OpenStudy (freckles):

awesome question and nice work on the explicit and non-explicit forms

OpenStudy (freckles):

this might be hard to show \[\lim_{n \rightarrow \infty}1 + \dfrac{1}{2+\dfrac{1}{2+\dfrac{1}{2+\cdots\text{(n times)}} }} =\sqrt{2}\]

OpenStudy (freckles):

I don't know enough about limits to show that is what I maen

OpenStudy (rational):

not realy that hard, at least not for you im pretty sure

OpenStudy (rational):

here is another interesting limit \[\lim_{n \rightarrow \infty}1 + \dfrac{1}{1+\dfrac{1}{1+\dfrac{1}{1+\cdots\text{(n times)}} }} =\phi\] \(\phi = \dfrac{\sqrt{5}+1}{2}\)

OpenStudy (rational):

turns out every irrational number has this infinite continued fraction form... so we can use this to approximate irrational numbers to any decimal place

OpenStudy (freckles):

kinda neat I think I remember the golden number being somehow related to Fibonacci sequence but I can't remember

OpenStudy (rational):

ratio of two consecutive fibonacci numbers approaches golden ratio as \(n\to \infty\)

OpenStudy (rational):

\[\lim\limits_{n\to \infty}\dfrac{F_{n+1}}{F_n} = \phi\] this one right ?

OpenStudy (freckles):

oh yeah you are right

OpenStudy (rational):

read about them in the morning haha yeah that limit is related to infinite continued fraction representaiton of \(\phi\)

OpenStudy (freckles):

According to this one site if I can somehow show that new ugly thing you wrote can come from \[\phi=1+\frac{1}{\phi} \\ \text{ then I can solve this for } \phi \\ \ \phi^2-\phi-1=0 \\ \phi=\frac{1 \pm \sqrt{1+4}}{2}=\frac{1 \pm \sqrt{5}}{2}\] when we can show that our sequence is increasing and not negative so we can throw out the negative solution

OpenStudy (freckles):

I called it ugly not because you wrote it in a ugly fashion but because it a seriously ugly looking fraction :p

OpenStudy (rational):

That is it! the same trick works for \(\sqrt{2}\) too !

OpenStudy (rational):

lol the word "ugly" is well taken :)

OpenStudy (freckles):

great news